OpenStudy (adamaero):

MEDAL ~ I'm trying to solve for v (not c;) .694c-c*c = v(c-v)

OpenStudy (adamaero):

I'm estimating it should be .75*3*10^8...but maybe I screwed up earlier in the problem

myininaya (myininaya):

what happened to c?

myininaya (myininaya):

\[.694c-c \cdot c=v(c-v)\] this is correct?

OpenStudy (adamaero):

I screwed something up before that. Nvr mind. Thanks tho